A 7.5 kg mass has its velocity change from 10.5 m/s east to 4.2 m/s east. Find the object's change in momentum.
1
Expert's answer
2020-11-02T15:07:49-0500

Solution. The change in momentum of an object is its mass times the change in its velocity

"\Delta p=m(v_f-v_i)"

where m=7.5kg is mass object; vf=4.2 m/s is final velocity; vi=10.5 m/s is initial velocity.

Let's choose east as a positive direction. As result


"\Delta p=7.5 kg (4.2 - 10.5)m/s=-47.25 kg\times m/s"

Answer. - 47.25 kg m/s (minus sign means decreasing momentum).